150.00 mL of 1.50 mol/L methanoic acid is being titrated with 0.750 mol/L sodium hydroxide. What is the pH after 30.00 mL of base has been added?

(Ka=1.78 x 10−4)
1
Expert's answer
2020-09-14T08:18:27-0400

Step I: Write balanced equation for the reaction,


HCOOH(aq) + NaOH(aq) "\to" HCOONa(aq) + H2O(l)


Mole ratio is 1:1


Initial Moles of HCOOH are calculated as;

1.5 moles are in 1000mL

? moles will be in 150mL

"=\dfrac{1.5mol x150mL}{1000mL}"


"=0.225 moles"


Moles of NaOH added is calculated as;


0.75 moles are in 1000mL

? moles will be in 30mL

"=\dfrac{0.75mol x 30mL}{1000mL}"


"=0.0225moles"


Therefore moles of HCOOH neutralized are = 0.0225 moles


And Moles of HCOOH remaining will be;


0.225-0.0225 = 0.02025 moles.


Step II: Concentration of acid and salt in the mixture


Total volume of solution becomes

150+30 = 180mL = 180/1000L =0.18L

Concentration of HCOOH remaining in the solution is calculated as;


"Concentration=\dfrac{number of moles}{Volume of solution in L}"


"Concentration = \dfrac{0.2025mol}{0.18L}"


= 1.125M


And Concentration of the salt (CHOONa) is;


"Concentration = \dfrac{0.0225mol}{0.18L}"


= 0.125M


Step III: Calculating pH


"pH=pKa + log\dfrac{[HCOONa]}{[HCOOH]}"



"pKa = -logKa = -log(1.78x10^-4) =3.75"


"\therefore pH = 3.75 + log\dfrac{0.125}{1.125}"


"pH = 3.75 + (-0.95) = 2.79"


pH of the solution is approximately 3